# On the correlation between Architectural Smells and Static Analysis Warnings

### v1

@@ 5 Discussion
This section discusses the results and presents lessons learned based on them.
Our findings reveal a moderate correlation between SAWs and ASs. However, different combinations of SATs and SAWs significantly influence the occurrence of ASs. Thus, our study’s key takeaway is that SAWs have a statistically discernible influence on ASs, despite the moderate correlation.
Furthermore, specific ASs show varying degrees of susceptibility to SAWs. For instance, as illustrated in Figure 4, UD’s acceptance or rejection of $H_{03}$ was almost equally probable. Likewise, SAWs that consistently fail to trigger ASs, i.e., Healthy, consistently reject the null hypothesis. As a result, this SAWs, acting as “healthy carriers”, can be safely discarded when prioritizing AS remediation efforts.
More specifically, the prevalence of FindBugs SAWs among the top co-occurring SAWs indicates that FindBugs might be more effective at finding issues related to ASs. This shows that FindBugs is effective at finding certain AS patterns. Thus, future research efforts should focus on the difference between FindBugs SAW and the other SATs SAWs. Thus grasping the characteristics that make FindBugs SAWs co-occur with all three ASs and with the “healthy carriers”.
Futhermore, future research efforts should focus on identifying UD AS, being the least co-occurring AS may suggest that these issues are not sufficiently characterized by the considered SAWs and SATs.
Therefore, by avoiding wasting remediation time on healthy carries, regarding $RQ_{3}$, our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in similar rankings to the optimal one. Healthy carriers represent 33.79% of our dataset. Hence we can safely drop a third of the SAW, focusing on the most AS-prone to effectively address AS without prior AS-severity or discovery knowledge.
@@ 6 Threats to Validity
Our specific design choices may impact our results, including our measurement process and data filtering. To address this threat, we based our choice on past studies [5, 22, 60, 46] and used well-established guidelines in designing our methodology [61, 40]. More specifically, we used four static analysis tools considered to be the most adopted by developers to detect software quality warnings [13] and ARCAN to detect the four architectural smells.
@@
Nevertheless, we acknowledge that the analyzed SAW does not represent all the possible SAWs currently available in the state-of-the-art. We mitigated this threat by selecting the most popular SAT. Nonetheless, future works will focus on extending the current selection of SAW. Similarly, we analyzed only three AS due to the lack of AS detection tools. Future works will focus on expanding the AS detectable by ARCAN to mitigate this threat.
@@
Internal Validity. Internal Validity is the extent to which an experimental design accurately identifies a cause-and-effect relationship between variables [39]. Our study relies on a large-scale analysis of 103 popular Java projects in the Qualitas Corpus dataset. However, they can potentially be biased from their project selection. We address this threat by designing our inclusion criteria based on past studies [60, 47, 22].
@@
External Validity. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [39]. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [39]. The analyzed project comes from the GitHub public repository. Mining versioning systems, particularly GitHub, also threaten external validity. More specifically, GitHub’s user base predominantly comprises developers and contributors to open-source projects, potentially skewing findings towards this specific demographic. In our context, this is not a real threat because it is the study’s focus. Moreover, the dynamic nature of GitHub, with frequent updates, forks, and merges, poses challenges in ensuring the stability and consistency of data over time. We addressed this issue by providing the raw data in our replication package. Furthermore, the accessibility of GitHub data is subject to various permissions and restrictions set by project owners, potentially hindering reproducibility and transparency in research. We addressed this research, limiting our data collection to public repositories. Nonetheless, the vast diversity of projects analyzed in our study presents a significant diversity in terms of programming languages, project sizes, and development methodologies, thus aiding the generalizability of results.
@@
Conclusion Validity focuses on how we draw conclusions based on the design of the case study, methodology, and observed results [39]. The distribution of our data determines the type of statistical tools we can use to test our hypothesis. We tested the data to asses if it departed from normality with the AD test, and we could reject the null hypothesis for each test. Therefore we selected the WT test instead of the t-test as Spearman’s $\rho$ in place of Pearsons. Furthermore, statistical tests threaten the conclusion’s validity regarding the appropriateness of statistical tests and procedures, such as assumption violation, multiple comparisons, and Type I or Type II errors. We address this issue using WT instead of the t-test due to the rejection of the normal data distributions. Moreover, we applied the HB correction [62] to balance type 1 and type 2 errors well. We also set our alpha to 0.01, hence reducing it from the standard value of 0.05 due to the many statistical tests we performed.
@@
Spearman’s $\rho$ is a non-parametric measure of the strength and direction of a correlation between two variables, but due to multiple factors, this approach may hinder the conclusion’s validity. For instance, the correlation coefficient is unreliable with a small sample size and cannot accurately represent the relationship between variables. Moreover, Spearman assumes a linear monotonic correlation between the variables; if the assumption were to prove invalid or not applicable, Spearman’s correlation might not accurately measure the strength and direction of the association. We have mitigated this issue by ensuring a representative sample and checking the data distribution. We could mitigate this issue by using other non-parametric measures of correlation. Still, our analysis found that Spearman’s $\rho$ was more suitable than other correlation measures, such as Kendall’s $\tau$, because it can effectively handle tied observations in the data. On the other hand, Kendall’s $\tau$ relies on concordant and discordant pairs of observations.
@@ 7 Related Work
This section discusses works related to our study regarding SATs and AS.
@@ 7.1 Static Analysis Tools
In this section, we report the relevant work on SAT focusing on their usage, warnings, and the detected issues [63, 64, 65]. SATs increasing in popularity [11, 66] given their ease-to-use [24]. Recently researchers focus on analyzing SAT usage  [67, 68, 69, 70] and effectiveness [13]. For instance, SATs effectively improve bug prediction models [71]. Developers use SAT to inspect software codebases, identify bugs [72, 73], evaluate code quality [74, 75, 76], and address coding issues [77, 78, 15].
Recent studies has explored SAT’s application across various programming languages [65], their configurations [65], and their evolution within software systems [79, 80, 77]. However, fewer studies have focused on the impact of SAT on software quality [79, 81, 15, 82, 83].
@@
Researchers have extensively studied SonarQube regarding the impact of SAW on software quality. They investigated the fault-proneness of SonarQube rules [79, 15, 88] and their susceptibility to changes [81, 15]. Findings indicate that SonarQube rules increase change-proneness at the class level, while correlations between faults and rules can reduce fault-proneness [79, 15].
Similarly, researchers have examined PMD warnings concerning their impact on pull request acceptance. Results show quality warnings do not affect pull request acceptance [82]. A comprehensive study [83] on warning trends in OSS projects using PMD highlighted that large-scale changes in SAT warnings often result from coding style modifications, such as braces and naming conventions. The presence of PMD in build configurations positively impacts defect density, suggesting an improvement in external quality.
Recently, Wright et al. [89] investigated why developers use SAT, revealing new tool requirements such as systems for recommending warnings based on developer expertise and collaborative interfaces for warning analysis. Conversely, Johnson et al. [74] examined why developers do not widely use SAT tools, identifying false positives and the presentation of warnings as significant barriers. Furthermore, Yang et al. [90] explored enhancing the utility of static code warning tools through data mining algorithms that filter out commonly ignored warnings, using FindBugs as a case study.
@@ 7.2 Architectural Smells
AS have been less thoroughly investigated than code smells. While numerous studies have explored correlations between code smells, research into the relationship between code smells and AS has been limited. One study found a minimal correlation between AS and code smells, suggesting that AS cannot be reliably inferred from code smells [19].
The connection between code anomalies, similar to code smells, and AS has also been studied with mixed results. Some findings indicated that many detected code anomalies did not correspond to architectural issues [17], while other results showed that over 80% of architectural problems were linked to code smells [18]. Specific code smells like Long Method, and God Class were consistently associated with architectural problems.
Oizumi et al. [91] investigate the aggregation of code anomalies and their relationship to AS and show that a single anomaly can signal an architectural problem. Specifically, 50% of syntactic aggregations and 80% of semantic aggregations are linked to design issues [92, 91]. A study does not find a strong correlation between AS and architectural degradation, outlining that AS cannot explain architectural degradation alone [93], while Sas et al. [94] highlights the effects of AS on long-term system maintainability and evolvability, supported by interview excerpts. Practitioners highlight which parts of the AS analysis offer actionable insights for planning refactoring activities.
@@
Finally, Fontana et al. [19] on the correlation between technical debt indices based on code-level issues from SonarQube revealed no correlation, suggesting differing impacts based on the features considered [26]. Although several tools have been developed for detecting architectural debt, including DV8, Designite, Jarchitect, and others, which focus on AS in various programming languages [96, 97, 98], no prior studies have examined correlations between AS and SAW as we have addressed in this paper.

### v2

@@ 5 Discussion
This section reflects on our findings, their practical implications, and how they compare to existing literature.
[…]
Our analysis revealed a moderate but statistically significant correlation between warnings and smells. While not strong enough to determine causation, the presence of certain warnings consistently aligned with specific smells across multiple systems. Interestingly, the strength of this correlation varied depending on the SAT and the warning type. FindBugs, in particular, produced a disproportionately high number of warnings that co-occurred with smells, suggesting that its rule set is more sensitive to architectural-level concerns.
Among the smells we studied, UD showed the least co-occurrence with warnings. This suggests that UD issues may lie outside the detection capabilities of standard SA tools or may manifest in ways not captured by low-level violations. Future research should explore alternative indicators or more abstract static metrics for such architectural erosion patterns [52, 8].
[…]
Interestingly, a handful of static analysis warnings surfaced as canaries in the coal mine, i.e., co-occurring with all three architectural smells (UD, HL, and CD). One such case is DMI_ EMPTY_ DB_ PASSWORD, which flags hardcoded empty database credentials. Beyond being a glaring security flaw, this warning hints at systemic issues: poor layering, tangled dependencies, and low cohesion. Another is EQ_ COMPARING_ CLASS_ NAMES, where equality checks rely on class names rather than proper instance comparisons—often a sign of brittle design and misuse of reflection or custom loading logic. Finally, DMI_ ENTRY_ SETS_ MAY_ REUSE_ ENTRY_ OBJECTS may appear trivial at first glance, but its misuse of shared objects can break encapsulation and introduce tight coupling across layers. While these individual violations may seem minor, their alignment with architectural smells hints they often live in “bad neighborhoods” of the codebase, regions suffering from architectural erosion on multiple fronts. Spotting these outliers early could help teams prevent cascading issues and prioritize fixes in areas that are both structurally weak and operationally risky.
[…]
Conversely, a large subset of warnings, 33.79% of our dataset, never co-occurred with any smells. We refer to these smells “non-co-occurring”. Since these warnings showed no statistical link to architectural erosion, they can be deprioritized when teams are constrained by time or resources. By focusing on the smell-prone subset of warnings, developers can streamline quality assurance efforts and reduce refactoring overhead.
[…]
Anyhow, we could question whether the observed co-occurrence between warnings and smells is genuinely informative, or merely a consequence of the high volume of warnings emitted by SATs. After all, SATs are known to generate alerts across most code entities. While this observation is valid, our contribution lies in distinguishing signal from noise.
Consider a common real-world scenario. A team working on an extensive legacy system integrates SonarQube, FindBugs, and PMD into their CI/CD pipeline. Over time, they are flooded with warnings, naming issues, unused variables, null checks, and developers begin to ignore them. Now imagine this team learns that specific warnings, such as smells NS_NON_SHORT_CIRCUIT or DM_GC, appear in over 75% of cases where architectural smells like CD are also present. This reframes those warnings from low-priority noise to meaningful indicators of deeper architectural problems.
Therefore, while warnings may be unsurprising, the statistical associations we uncover, based on structured co-occurrence patterns, hypothesis testing, and severity-aware prioritization, provide valuable insights. We demonstrate that not all warnings are equally relevant: some consistently co-occur with smells and serve smells lightweight architectural risk indicators, while others (”non-co-occurring smells”) do not. This finding enables developers to focus their efforts more effectively, particularly in fast-paced settings where architectural tools like ARCAN are unavailable or impractical.
[…]
5.1 Previous Study Results
From a practical perspective, these findings offer concrete guidance. Development teams can confidently filter out non-co-occurring smells and instead concentrate on warnings that frequently co-occur with smells. When explicit smell severity data is unavailable, they can prioritize remediation based on either warning severity levels or warning–smell co-occurrence probabilities. Both strategies, as shown in our ranking analysis (Figure 5), approximate the optimal remediation order based on smell severity.
[…]
For example, SonarQube users could enrich their dashboards with co-occurrence probabilities to better triage warnings. At the same time, reviewers could flag high-risk FindBugs warnings, such as smells that are frequently linked to CD, and smells candidates for early intervention. Although this method does not replace architecture-specific analysis tools, it provides an accessible, lightweight alternative that integrates seamlessly into existing CI/CD workflows.
[…]
These insights directly address $RQ_{3}$, which explores how warning-based indicators can support smells remediation. Our prioritization strategies, particularly those using warning severity and co-occurrence likelihood, produced rankings that closely matched smells-severity-based rankings. This shows that architectural concerns can be anticipated, at least in part, through patterns already embedded in developer workflows.
Compared to prior work such smells Fontana et al. [28], which found weak correlations between code smells and smells, our study reveals that certain warnings, though more granular and often disregarded, carry meaningful signals about architectural issues. This distinction highlights our novel contribution: bridging SAT output with architectural-level insights, while maintaining compatibility with tools developers already use.
Our findings also align with those of Macia et al. [54], who showed that code anomalies introduced early in a project often lead to architectural erosion. Their work emphasized manually validated anomalies in a limited set of systems. By contrast, our study generalizes this observation across 103 projects, using automatically extracted warnings from widely adopted tools. warnings offer a more representative and fine-grained lens into the types of quality issues developers regularly encounter.
We also extend the insights of Oizumi et al. [68], who found that semantic aggregations of code anomalies better indicate architectural problems than individual instances. Our study confirms that aggregated co-occurrence patterns, especially from tools like FindBugs, are useful in estimating the likelihood and severity of smells. Moreover, we provide a prioritization framework grounded in these co-occurrence metrics, offering a more scalable and automation-friendly solution.
5.2 Implications for Researchers.
This study opens up several research directions. First, while we confirmed a moderate correlation between warnings and smells, the variation across tools and smells suggests that deeper semantic properties influence these links. Future work could explore contextual factors such smells code structure, change history, or architectural annotations to improve prediction models.
Second, the concept of “non-co-occurring smells” introduces a novel dimension for understanding warning reliability. Investigating what makes these warnings architecturally benign, e.g., warning type, location, or context, could help refine rule sets for static analysis tools, minimizing false positives and increasing developer trust.
Finally, given the scalability of our approach across 72 million lines of code and over 800,000 classes, future studies can explore transfer learning or AI-driven prioritization across repositories and ecosystems. This could lead to generalizable models for quality-aware development that adapt to project characteristics and team practices.
[…]
Overall, our work serves smells a bridge between static analysis and architectural health monitoring, bringing together low-level tooling and high-level design quality in a way that supports both daily development and long-term maintainability.
Our finding also opens the door to a new class of tooling: one that connects co-occurrence signals across multiple warnings and smells to detect emergent architectural pathologies. For instance, a cluster of warnings repeatedly linked to UD, HL, and CD could indicate a nascent God Component, Blob, or Lava Flow, patterns that rarely surface through individual metrics alone. By layering co-occurrence with topology- or dependency-based metrics, future tools could provide early, lightweight detection of complex architectural anti-patterns. This would allow teams to intervene before such structures fully take shape and become costly to disentangle.
[…]
@@ 6 Threats to Validity
Our specific design choices may impact our results, including our measurement process and data filtering. To address this threat, we based our selection on past studies [3, 90, 81] and used well-established guidelines in designing our methodology [76, 7]. More specifically, we used three static analysis tools considered to be the most adopted by developers to detect software quality warnings [24] and ARCAN to detect the four architectural smells.
@@
Nevertheless, we acknowledge that the analyzed warning does not represent all the possible warnings currently available in the state-of-the-art. Our study selected the most popular SATs. Nonetheless, future works will focus on extending the current selection of warnings. Similarly, we analyzed only three AS due to the lack of AS detection tools.
Moreover, the original data is not recent. Therefore, it can be argued that it may face deprecation. Similar arguments can be put towards the SAT. It is worth noticing that deprecation, in this context, is a multifaceted threat. Firstly, according to [22], deprecation cannot impact generalizability. Deprecation, in this specific context, may affect the quantity of AS detected, given that the concept of microservices has recently reached its peak of interest [^7] https://content.ardoq.com/gartner-magic-quadrant-for-enterprise-architecture-tools.
@@
Internal Validity. Internal validity is the extent to which an experimental design accurately identifies a cause-and-effect relationship between variables [100]. Moreover, according to  [100], internal validity not only concerns causal relationships, which our study does not directly claim, but also involves controlling for bias or confounding factors, ensuring consistency in data collection, and avoiding measurement errors. Indeed, our study relies on a large-scale analysis of 103 popular Java projects in the Qualitas Corpus dataset. However, they can potentially be biased by their project selection. We address this threat by designing our inclusion criteria based on past studies [91, 90]. Futhermore, we addressed tool selection bias by including multiple SATs widely adopted in industry and using ARCAN, one of the few validated tools for AS detection. Our findings remain valid within the scope of the selected tools.
@@
External Validity. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [100]. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [100]. Our study is based on the Qualitas Corpus dataset, which was last updated in 2013. As such, it includes monolithic Java systems and does not reflect recent architectural paradigms, such as microservices or cloud-native systems. This limits the generalizability of our findings to modern software projects. The dynamics between warnings and smells in contemporary ecosystems may differ due to design practices, modularity, deployment infrastructure, and static analysis tooling changes. Future work should replicate this study on newer datasets reflecting current architectural trends. The analyzed project comes from the GitHub public repository. Mining versioning systems, particularly GitHub, also threaten validity. More specifically, GitHub’s user base predominantly comprises developers and contributors to open-source projects, potentially skewing findings towards this specific demographic. In our context, this is not a real threat because it is the study’s focus. Moreover, the dynamic nature of GitHub, with frequent updates, forks, and merges, poses challenges in ensuring the stability and consistency of data over time. We addressed this issue by providing the raw data in our replication package. Furthermore, the accessibility of GitHub data is subject to various permissions and restrictions set by project owners, potentially hindering reproducibility and transparency in research. We addressed this research, limiting our data collection to public repositories. Regarding generalizability, the vast diversity of projects analyzed in our study presents a significant diversity in terms of programming languages, project sizes, and development methodologies, thus aiding the generalizability of results.
@@
Conclusion Validity focuses on how we draw conclusions based on the design of the case study, methodology, and observed results [100]. The distribution of our data determines the type of statistical tools we can use to test our hypothesis. We tested the data to asses if it departed from normality with the AD test, and we could reject the null hypothesis for each test. Therefore, we selected the WT test instead of the t-test as Spearman’s $\rho$ in place of Pearson’s. Furthermore, statistical tests threaten the validity of the appropriateness of statistical tests and procedures, such as assumption violation, multiple comparisons, and Type I or Type II errors. We address this issue using WT instead of the t-test due to the rejection of the normal data distributions. Moreover, we applied the HB correction [41] to balance type 1 and type 2 errors well. We also set our alpha to 0.01, hence reducing it from the standard value of 0.05 due to the many statistical tests we performed.
@@
Spearman’s $\rho$ is a non-parametric measure of the strength and direction of a correlation between two variables, but due to multiple factors, this approach may hinder the validity. For instance, the correlation coefficient is unreliable with a small sample size and cannot accurately represent the relationship between variables. Moreover, Spearman assumes a linear monotonic correlation between the variables; if the assumption were to prove invalid or not applicable, Spearman’s correlation might not accurately measure the strength and direction of the association. We have mitigated this issue by ensuring a representative sample and checking the data distribution. We could reduce this issue by using other non-parametric measures of correlation. Still, our analysis found that Spearman’s $\rho$ was more suitable than other correlation measures, such as Kendall’s $\tau$, because it can effectively handle tied observations in the data. On the other hand, Kendall’s $\tau$ relies on concordant and discordant pairs of observations.
@@ 7 Related Work
This section discusses works related to our study regarding SATs and smells. Table 8 compares our work with prior studies on static analysis tools and architectural smells. Existing research has investigated SAT usage, developer perceptions, or the evolution of smells, yet none examined direct correlations between SAT warnings and architectural smells. According to Table 8, our study distinguishes itself by providing the first large-scale empirical evidence on their relationship and by proposing warning-based prioritization strategies that effectively approximate remediation guided by architectural impact severity. Therefore, we position our contribution as a pragmatic bridge between ubiquitous SAT outputs and higher-level architectural erosion.
Table 8: Comparison of our work with related studies on SATs and smells.
Study Focus Method / Scope Contributions and Gaps
Falessi et al. [25], Lenarduzzi et al. [48], Tollin et al. [92] SAT rules and quality impact Empirical studies on SonarQube rules, class-level correlations Showed links between warnings and change/fault-proneness, but did not study architectural smells
Trautsch et al. [94], Lenarduzzi et al. [49] Warning evolution and trends Longitudinal analyses with PMD, PR acceptance Studied evolution and acceptance, but no connection with architectural issues
Johnson et al. [38], Wright et al. [101] Developer perceptions of SATs Interviews, surveys, user studies Identified adoption barriers (e.g., false positives), but no quantitative correlation with architecture
Yang et al. [103], Ge et al. [32], Li and Yang [50] Improving SAT utility Data mining, preference mining, automated tracking Enhance prioritization of warnings, but focus on code-level issues
Fontana et al. [28], Macia et al. [54], Oizumi et al. [68], Herold [35], Sas et al. [82] Architectural smells and erosion Correlation studies, anomaly aggregation, longitudinal analyses Explored code–architecture links, but did not involve SAT warnings
Gnoyke et al. [33], Jolak et al. [39] Smell evolution and tooling Novel tools and empirical studies Extend smell detection, but do not combine SATs with smells
Simões and Venson [84], Yıldırım et al. [104], Nizam et al. [65] LLMs as static analysis Benchmarks vs. SATs, embedding optimization Emerging use of AI for quality checks, no empirical correlation with architectural smells
Our work SAT warnings $\leftrightarrow$ architectural smells Large-scale study (103 projects, 72M LOC), 3 SATs + ARCAN First to reveal moderate correlations, propose prioritization strategies bridging code-level warnings and architectural remediation
@@ 7.1 Static Analysis Tools
In this section, we report the relevant work on SAT focusing on their usage, warnings, and the detected issues [26, 34, 9]. In recent years, SATs have strongly increased in popularity [96, 45] given their ease-to-use [106]. Along with this attention increase, researchers have focused on analyzing SAT usage  [97, 63, 107, 64] and effectiveness [24]. For instance, SATs effectively improve bug prediction models [73]. Developers use SAT to inspect software codebases, identify bugs [77, 93], evaluate code quality [38, 83, 55], and address coding issues [79, 43, 48].
Recent studies have explored SAT’s application across various programming languages [9], their configurations [9], and their evolution within software systems [25, 18, 79]. However, fewer studies have focused on the impact of SAT on software quality [25, 92, 48, 49, 94].
@@
Researchers have extensively studied SonarQube regarding the impact of warning on software quality. They investigated the fault-proneness of SonarQube rules [25, 48, 46] and their susceptibility to changes [92, 48]. Findings indicate that SonarQube rules increase change-proneness at the class level, while correlations between faults and rules can reduce fault-proneness [25, 48].
Similarly, researchers have examined PMD warnings concerning their impact on pull request acceptance. Results show quality warnings do not affect pull request acceptance [49]. A comprehensive study [94] on warning trends in OSS projects using PMD highlighted that large-scale changes in SAT warnings often result from coding style modifications, such smells braces and naming conventions. The presence of PMD in build configurations positively impacts defect density, suggesting an improvement in external quality.
Recently, Wright et al. [101] investigated why developers use SAT, revealing new tool requirements such smells systems for recommending warnings based on developer expertise and collaborative interfaces for warning analysis. Conversely, Johnson et al. [38] examined why developers do not widely use SAT tools, identifying false positives and the presentation of warnings smells significant barriers. Furthermore, Yang et al. [103] explored enhancing the utility of static code warning tools through data mining algorithms that filter out commonly ignored warnings, using FindBugs smells in a case study. Similarly, Ge et al. [32] proposed preference mining to recommend the most suitable SAT results for each project accordingly. Li and Yang [50] studied the benefits of automating the tracking of the evolution of static code warnings with StaticTracker, a tool that combines results from PMD and SpotBugs, successor of FindBugs, for static bug detection.
Along with the rise of new techniques such as Large Language Models (LLMs), Simões and Venson [84] evaluated the overall quality of code with LLMs as SATs, and compared their results with those from SonarQube. Yıldırım et al. [104] scaled up this research trend by comparing 10 SATs against four LLMs on assessing their ability to detect a compilation of 40 API vulnerabilities in the source code. Furthermore, Nizam et al. [65] explored different embedding techniques to improve the quality of code embeddings when leveraging LLMs for source code analysis.
@@ 7.2 Architectural Smells
Smells have been less thoroughly investigated than code smells. While numerous studies have explored correlations between code smells, research into the relationship between code smells and architectural smells has been limited. One study found a minimal correlation between smells and code smells, suggesting that smells cannot be reliably inferred from code smells [28].
The connection between code anomalies, similar to code smells, and smells has also been studied with mixed results. Some findings indicated that many detected code anomalies did not correspond to architectural issues [54], while other results showed that over 80% of architectural problems were linked to code smells [53]. Specific code smells like Long Method, and God Class were consistently associated with architectural problems.
Oizumi et al. [68] investigate the aggregation of code anomalies and their relationship to smells and show that a single anomaly can signal an architectural problem. Specifically, 50% of syntactic aggregations and 80% of semantic aggregations are linked to design issues [67, 68]. A study does not find a strong correlation between smells and architectural erosion, outlining that smells cannot explain architectural erosion alone [35], while Sas et al. [82] highlights the effects of smells on long-term system maintainability and evolvability, supported by interview excerpts. Practitioners highlight which parts of the smells analysis offer actionable insights for planning refactoring activities.
@@
Recently, Gnoyke et al. [33] studied the evolution patterns of smells to describe symptoms of architectural erosion. They present extended techniques to measure either intra-version as well as inter-version smells. Following the latest works on smells, Jolak et al. [39] introduced a novel open-source smell analysis tool, which gathers and analyzes data related to smells, modularity and testability. Additionally, they studied the statistical relationship between different types of smells with modularity and testability.
Finally, Fontana et al. [28] on the correlation between technical debt indices based on code-level issues from SonarQube revealed no correlation, suggesting differing impacts based on the features considered [75]. Although several tools have been developed for detecting architectural debt, including DV8, Designite, Jarchitect, and others, which focus on smells in various programming languages [80, 31, 11, 39], no prior studies have examined correlations between smells and warning smells we have addressed in this paper.